In the 2004 summer Olympics, France won more gold medals than Italy, who won more gold metals than Korea. If the total number of gold medals won by these three countries is three consecutive integers whose sum is 20, find the number of gold metals won by each.
Answer by checkley71(8403)

X+(X+1)+(X+2)=20 OR 3X-3=20 3X=20-3 OR 3X=17 OR X=17/3 X=5 2/3 THUS THE OTHER TWO NUMBERS ARE 6 2/3 & 7/2/3.
PROOF 5 2/3+6 2/3+7 2/3=20 OR 18 6/3 OR OR 20
THESE ARE NOT INTERGERS!